# Interpretable Vision-Language Survival Analysis with Ordinal Inductive Bias for Computational Pathology

### v3

@@ 3.3 Ordinal Incidence Function Prediction
The prediction target of VL-based models is class probability, usually calculated based on the similarity measures between image features and a group of textual class prompts. To make survival prediction compatible with this way, we propose to take individual incidence function as prediction target and introduce another ordinal inductive bias term into it, as shown in Figure 1(b).
@@
where $\tau$ is a temperature parameter optimized in training. From a conventional classification perspective, $\hat{y}_{c}$ given by Eq. (6) can be cast as the probability that an event first occurs at time $c$; $\hat{{\bm{y}}}$ is thus a probability distribution on the first hitting time. Such type of prediction is closely associated with the concept of incidence function (IF) (Fine & Gray, 1999) and the first hitting time model (Lee & Whitmore, 2006; Lee et al., 2018) in traditional SA. Therefore, we call Eq. (6) the prediction of individual IF and leverage IF tools for our task. By definition, cumulative IF is $\text{CIF}(c)=\sum_{i=1}^{c}\hat{y}_{c}$. Survival function (surviving past time $c$) can be written as $\hat{S}(c)=1-\text{CIF}(c)=1-\sum_{i=1}^{c}\hat{y}_{i}$.
Ordinal Inductive Bias in Incidence Function As the predictive probability $\hat{{\bm{y}}}$ is distributed over ordered survival classes, we consider an ordinality constraint for this distribution (see Appendix C for further elucidations). Concretely, given that $c$ is the survival class with the largest probability, we assume there is a consecutive decline in probability for those classes away from $c$, as shown in Figure 1(b). For this purpose, we adopt Earth Mover’s Distance (EMD) (Levina & Bickel, 2001) as the measure to quantify the distance between $\hat{{\bm{y}}}$ and ${\bm{y}}$ (ground truth distribution), written as
@@
where $\text{CDF}(\cdot)$ means cumulative distribution function. EMD is a measure aware of distribution ordinality. Please refer to Appendix B.3 for more details. Furthermore, we adopt a squared EMD objective (Hou et al., 2017) to regularize the predictive distribution $\hat{{\bm{y}}}$ in model optimization:
@@
We define ${\bm{y}}(c,\delta)$ as follows. For the patient with event, ${\bm{y}}(c,\delta=1)=\text{Softmax}\big{(}\tau^{\prime}\cdot(2\cdot{\mathbb{I}}_{i=c}-1)\big{)}$. For the patient censored at time $c$, we only know that its actual time-to-event is not less than $c$, so ${\bm{y}}(c,\delta=0)=\text{Softmax}\big{(}\tau^{\prime}\cdot(2\cdot{\mathbb{I}}_{i\geq c}-1)\big{)}$. ${\mathbb{I}}_{i}\in\{0,1\}^{C}$ is an indicator with element ‘1’ at index $i$. The value of $\tau^{\prime}$ is from $\tau$ but not involved in optimization. We note that the ordinality introduced to ${\bm{F}}_{\text{text}}$ and $\hat{{\bm{y}}}$ could be cast as model-level and subject-level inductive bias, respectively; still, there are intriguing connections between them. Detailed discussions could be found in Appendix C.
@@ 3.4 Overall Training Objectives
Furthermore, with the $\mathcal{L}_{\text{EMD}}$ given in Eq. 8, our overall training objective is minimizing
@@ 3.5 Prediction Interpretation
Understanding how a model makes predictions is crucial for reliable decision-making, especially in medical domain. We thereby study the prediction behavior of VLSA and propose an interpretation method based on the classic Shapley values (see Appendix B.4) from cooperative game theory.
Concretely, we interpret the survival prediction of VLSA through language-encoded prognostic priors, i.e., ${\bm{P}}=[{\bm{p}}_{1},\dots,{\bm{p}}_{M}]^{\mathsf{T}}$ described in Section 3.1, because i) each prior intuitively describes a certain prognostic visual feature in WSIs and ii) these language-encoded priors play an important role in learning valuable prognostic clues and making final survival predictions. However, each prior term’s contribution to risk prediction cannot be computed directly since it is not linearly correlated with $\hat{{\bm{y}}}$. Thus, based on a principled framework for model interpretation—Shapley value (Shapley, 1953; Lundberg, 2017), we calculate the contribution of each prior term by
@@
where ${\mathbb{P}}$ is a set of prognostic priors $\{{\bm{p}}_{1},\dots,{\bm{p}}_{M}\}$. The function $f_{\text{risk}}({\mathbb{Z}},{\bm{X}})$ outputs the risk prediction given prognostic prior subset ${\mathbb{Z}}$ and WSI features ${\bm{X}}$ as inputs. Specifically, we i) calculate ${\bm{f}}_{\text{image}}$ using a subset of ${\bm{F}}$ corresponding to ${\mathbb{Z}}$, ii) use it to obtain the IF prediction $\hat{{\bm{y}}}$ through Eq.(6), and iii) finally derive the risk via the summation of CIF: $\hat{R}=\sum_{i=1}^{C}\text{CIF}(i)$.
@@ 4.1 Experimental Setup
Datasets Following Chen et al. (2021), five publicly-available datasets from TCGA (The Cancer Genome Atlas) are used in experiments: BLCA (bladder urothelial carcinoma) ($N$ = 373), BRCA (breast invasive carcinoma) ($N$ = 956), GBMLGG (glioblastoma & lower grade glioma) ($N$ = 569), LUAD (lung adenocarcinoma) ($N$ = 453), and UCEC (uterine corpus endometrial carcinoma) ($N$ = 480). These datasets cover five different cancer types, a total number of 2,831 patients, and 3,530 diagnostic gigapixel WSIs. We adopt a standard CLAM tool (Lu et al., 2021) to preprocess all WSIs into multi-instance bags for training. CONCH (Lu et al., 2024), as a state-of-the-art foundational VLM recently developed for CPATH, is employed to provide powerful encoders $E_{\text{image}}$ and $E_{\text{text}}$. Refer to Appendix D.1 for more details on datasets.
Baselines Vision-only and vision-language methods are compared in experiments. Vision-only (V) baselines contain ABMIL (Ilse et al., 2018), TransMIL (Shao et al., 2021b), ILRA (Xiang & Zhang, 2023), DeepAttnMISL (Yao et al., 2020), and Patch-GCN (Chen et al., 2021). Vision-language (VL) baselines are MI-Zero (Lu et al., 2023), CoOp (Zhou et al., 2022), and OrdinalCLIP (Li et al., 2022). MI-Zero is a VL-based zero-shot approach for classification; we adapt it for SA to assess the lower bound performance of VLM on new SA tasks. As CoOp and OrdinalCLIP are not proposed for gigapixel WSIs, we adapt them for comparison: i) in vision-end, the frequently-adopted attention-based MIL, i.e., ABMIL, is used for instance aggregation; ii) in language-end, CoOp optimizes both context and class prompts without ordinal inductive bias, while OrdinalCLIP improves CoOp using ordinal class prompts. All methods use the same $E_{\text{image}}$ from CONCH; VL-based methods involve an additional VL projection layer. More details are provided in Appendix D.2 and D.4.
Evaluation Metrics To evaluate models more comprehensively, we follow traditional SA (Qi et al., 2023a; b) to consider multiple metrics. (1) Concordance index (CI), as a prevalent metric in SA, assesses models’ discrimination power. (2) Mean absolute error (MAE) gives the absolute error of model’s prediction on time-to-event. (3) Distribution calibration (D-cal(Haider et al., 2020) is a statistical test to evaluate models’ ability in calibrating survival distribution prediction; we mainly check if there is no significant difference in distribution between ground truth and prediction ($p>$ 0.05). For any method, we adopt 5-fold cross-validation to evaluate its performance and report the average on 5 folds. Please refer to Appendix D.3 for more details on evaluation metrics.
4.2 Comparison with Baselines
@@ 4.3 Analysis of Language-Encoded Prognostic Priors
Figure 2: Interpreting the survival prediction of VLSA via language-encoded prognostic priors. Top row shows language descriptions (simplified for better view) about prognostic visual features in WSIs. Detailed texts are provided in Appendix D.4. Middle row gives the most representative patches corresponding to each prognostic text. Last row presents each prognostic prior’s contribution to overall risk. We mainly examine the top three language priors in terms of contribution. The three examples are from the first three datasets. Refer to Appendix E for more results.
Visualization We conduct a case study to further examine the role of language-encoded prognostic priors in our VLSA. As shown in Figure 2, we observe that valuable visual features can be highlighted by prognostic texts in most cases. These visual features, e.g., “differentiation of tumor cells” and “high cellular density”, could often provide important clues for cancer prognosis. We note that, such desirable behavior is largely attributed to the intrinsic property of foundational VLMs, i.e., VLMs offer a well-aligned vision-language embedding space to enable us to manipulate visual features with language prior. We believe that this property is particularly important to the MIL in which a model is usually required to aggregate numerous instances under weak supervision.
@@ 4.4 Analysis of Ordinal Inductive Biases
Ablation Study We conduct an ablation study to verify the effectiveness of our two ordinal inductive bias terms. Its result is shown in Table 3. We have two main observations. (1) Ordinal survival prompt makes a large contribution (3.4% and 3.6%) to the VLSA’s performance on BRCA in terms of CI; While on the other four datasets, the contribution is often marginal. (2) Ordinal probability, i.e., ordinal IF, could often help to decrease MAE. More discussions on these two inductive bias terms are provided in Appendix C.
@@
Ordinality Visualization (1) Survival prompts: from Figure 3 (a), we can see that survival prompts are better in ordinality when considering an ordinal inductive bias for them. (2) Predictive incidence function: the case study shown in Figure 3 (b) suggest that the ordinality constraint imposed on IF prediction could help to decrease the EMD between $\hat{{\bm{y}}}$ and ground truth distribution. This leads to the improvements of MAE (see Table 3).
Figure 3: Ordinality visualization. (a) Survival prompts. Acc is the accuracy of prompt ranking. The results are from the first three datasets. More results are shown in Appendix E. (b) Predictive probability (Proba.) w/o and w/ ordinality. The patients from test set are used for prediction. $\Delta$EMD is equal to $\text{EMD}({\bm{y}},\hat{{\bm{y}}}_{\text{w/o}})-\text{EMD}({\bm{y}},\hat{{\bm{y}}}_{\text{w/}})$. Vertical dashed line indicates individual time-to-event.
4.5 Few-Shot Survival Prediction
We design few-shot learning (FSL) scenarios to further compare different SA methods. Since existing works on SA rarely study this, this experiment follows common few-shot settings. Concretely, we randomly sample $s$ patients from each time-discrete class. For censored patients, their true class labels are partially given, so we adopt the KM method (Kaplan & Meier, 1958) to estimate them in advance for few-shot sampling. Moreover, due to small patient numbers, we run each single experiment 5 times and report the median metrics, following CONCH (Lu et al., 2024).
Table 4: Few-shot survival prediction. The CI averaged on five datasets is presented.
@@ 4.5 Few-Shot Survival Prediction
The results of few-shot survival prediction are shown in Table 4. (1) From these results, we find that our VLSA still often obtains the best performance in FSL scenarios. This indicates the data efficiency of our VLSA. (2) When using only 16 shots, our VLSA can obtain comparable performance with nearly-all full-shot baseline models. (3) Furthermore, we observe that, in standard full-shot setting, two VL-based baselines, i.e., CoOp and OrdinalCLIP with ABMIL as their MIL aggregator, only show an improvement of 0.6% and 1.8% over vision-only ABMIL, respectively. By contrast, the average improvements seen in few-shot settings are around 6.5%, much larger than those in full-shot. This observation also suggests the advantage of VL-based schemes over vision-only ones in terms of data efficiency.
4.6 Analysis of Model Efficiency
We evaluate the efficiency of all models in this experiment. The number of learnable parameters (# Params) and Multiply–Accumulate operations (# MACs) are taken as main evaluation metrics. From the results shown in Table 5, we can see that VLSA not only has fewer learnable parameters (284.16K) than all baseline models except for ABMIL but also is the most efficient one (0.25G) in terms of model inference cost. Our further analysis is as follows.
Table 5: Model efficiency. A random WSI (from BLCA) with 20,265 instances is used to measure the # MACs in model inference.
Model # Params # MACs
V ABMIL 137.74K 2.68G
TransMIL 683.55K 14.23G
ILRA 3.16M 34.79G
DeepAttnMISL 329.22K 2.68G
Patch-GCN 363.91K 3.41G
VL CoOp 434.18K 2.68G
OrdinalCLIP 409.60K 2.68G
VLSA 284.16K 0.25G
5 Limitation and Future Work
Although promising results are obtained in this study, we note that there are still some limitations. (1) The number of datasets and the diversity of cancer types are limited. Due to the extremely-high resolution of histopathology WSIs, it is usually challenging to collect tens of thousands of images and preprocess them into feasible training data. (2) The textual descriptions about cancer-specific prognostic prior are derived from GPT-4o. They may be lacking in knowledge completeness and accuracy. More holistic and accurate prognostic prior knowledge are likely to provide stronger guidance for weakly-supervised MIL and perform better in survival prediction. (3) The foundational VLM that our empirical results rely on is CONCH (Lu et al., 2024), a recent state-of-the-art VLM for CPATH. More pathology VLMs may be needed to further validate our VLSA.

### v4

@@ 3.3 Ordinal Incidence Function Prediction
The prediction target of VL-based models is usually class probability, calculated based on the similarity measures between image features and a group of textual class prompts. To make survival prediction compatible with this way, we propose to take individual incidence function as prediction target and introduce another ordinal inductive bias term into it, as shown in Figure 1(c).
@@
where $\tau$ is a temperature parameter optimized in training. From a conventional classification perspective, $\hat{y}_{c}$ given by Eq. (6) can be cast as the probability that an event first occurs at time $c$; $\hat{{\bm{y}}}$ is thus a probability distribution on the first hitting time. Such prediction is closely associated with the concept of incidence function (IF) (Fine & Gray, 1999) and the first hitting time model (Lee & Whitmore, 2006; Lee et al., 2018) in traditional SA. Therefore, we call Eq. (6) the prediction of individual IF and leverage IF tools for the interested task. By definition, cumulative IF is $\text{CIF}(c)=\sum_{i=1}^{c}\hat{y}_{c}$. Survival function (surviving past time $c$) can be written as $\hat{S}(c)=1-\text{CIF}(c)=1-\sum_{i=1}^{c}\hat{y}_{i}$.
Ordinal Inductive Bias in Incidence Function As $\hat{{\bm{y}}}$ is a probability distributed over ordered survival classes, we consider an ordinality constraint for it (see Appendix B for further elucidations). Concretely, given that $c$ is the survival class with the largest probability, we assume there is a consecutive decline in probability for those classes away from $c$, as shown in Figure 1(c). Note that our assumption is concerned with the event that first occurs, as $\hat{{\bm{y}}}$ is defined as the probability distribution on the first hitting time. To impose the ordinality constraint on $\hat{{\bm{y}}}$, at first we adopt Earth Mover’s Distance (EMD) (Levina & Bickel, 2001) as the measure to quantify the distance between $\hat{{\bm{y}}}$ and ${\bm{y}}$ (ground truth distribution), written as
@@
where $\text{CDF}(\cdot)$ means cumulative distribution function. EMD is a measure aware of distribution ordinality as it considers the geometry property of distribution in distance measurement and it is smaller when the geometry (shape) of two distributions is closer. Please refer to Appendix A.3 for detailed explanations. Furthermore, we adopt a squared EMD objective (Hou et al., 2017) to regularize $\hat{{\bm{y}}}$ in model optimization:
@@
${\bm{y}}(c,\delta)$ is defined as follows. For the patient with event, ${\bm{y}}(c,\delta=1)=\text{Softmax}\big{(}\tau^{\prime}\cdot(2\cdot{\mathbb{I}}_{i=c}-1)\big{)}$. For the patient censored at time $c$, we only know that its actual time-to-event is not less than $c$, so ${\bm{y}}(c,\delta=0)=\text{Softmax}\big{(}\tau^{\prime}\cdot(2\cdot{\mathbb{I}}_{i\geq c}-1)\big{)}$. ${\mathbb{I}}_{i}\in\{0,1\}^{C}$ is an indicator with element ‘1’ at index $i$. The value of $\tau^{\prime}$ is from $\tau$ but not involved in optimization. We further analysis and discuss the ordinal inductive biases introduced into ${\bm{F}}_{\text{text}}$ and $\hat{{\bm{y}}}$; the details can be found in Appendix B.
@@ 3.4 Overall Training Objectives
For censored patients ($\delta=0$), $\mathcal{L}_{\text{MLE}}$ minimizes $\sum_{i=1}^{c-1}\hat{y}_{i}$, i.e., the probability that the event of interest first occurs at discrete bins 1, 2, $\dots$, or $c-1$, according to the definition of $\hat{y}_{i}$ in Eq. (6). Furthermore, with the $\mathcal{L}_{\text{EMD}}$ given in Eq. (8), our overall training objective is minimizing
@@ 3.5 Prediction Interpretation
Understanding how a model makes predictions is crucial for reliable decision-making, especially in medical domains. We thereby study the prediction behavior of VLSA and propose an interpretation method based on the classic Shapley values (Appendix A.4) from cooperative game theory.
Concretely, we interpret the survival prediction of VLSA through language-encoded prognostic priors, i.e., ${\bm{P}}=[{\bm{p}}_{1},\dots,{\bm{p}}_{M}]^{\mathsf{T}}$ described in Section 3.1, because i) each prior intuitively describes a certain prognostic visual feature in WSIs and ii) these language-encoded priors play an important role in learning valuable prognostic clues and making final survival predictions. However, each prior item’s contribution to risk prediction cannot be computed directly since it is not linearly correlated with $\hat{{\bm{y}}}$. Thus, based on a principled framework for model interpretation—Shapley value (Shapley, 1953; Lundberg, 2017), we calculate the contribution of each prior item by
@@
where ${\mathbb{P}}$ is a set of prognostic priors $\{{\bm{p}}_{1},\dots,{\bm{p}}_{M}\}$. The function $f_{\text{risk}}({\mathbb{Z}},{\bm{X}})$ outputs the risk prediction given the prognostic prior subset ${\mathbb{Z}}$ and the WSI features ${\bm{X}}$ as inputs. Specifically, we i) calculate ${\bm{f}}_{\text{image}}$ using a subset of ${\bm{F}}$ corresponding to ${\mathbb{Z}}$, ii) use it to obtain the IF prediction $\hat{{\bm{y}}}$ through Eq. (6), and iii) finally derive the risk via the summation of CIF: $\hat{R}=\sum_{i=1}^{C}\text{CIF}(i)$.
@@ 4.1 Experimental Setup
Datasets Following Chen et al. (2021a), five publicly-available datasets from TCGA (The Cancer Genome Atlas) are used in experiments: BLCA (bladder urothelial carcinoma) ($N$ = 373), BRCA (breast invasive carcinoma) ($N$ = 956), GBMLGG (glioblastoma & lower grade glioma) ($N$ = 569), LUAD (lung adenocarcinoma) ($N$ = 453), and UCEC (uterine corpus endometrial carcinoma) ($N$ = 480). These datasets cover five different cancer types, a total number of 2,831 patients, and 3,530 diagnostic gigapixel WSIs. Overall Survival (OS), which only occurs once, is set as the event of interest in this paper, following Chen et al. (2021a). We adopt a standard CLAM tool (Lu et al., 2021) to preprocess all WSIs into multi-instance bags for training. CONCH (Lu et al., 2024), as a state-of-the-art VLM recently developed for CPATH, is utilized to provide powerful encoders $E_{\text{image}}$ and $E_{\text{text}}$. More details on datasets can be found in Appendix C.1.
@@
Evaluation Metrics To evaluate models more comprehensively, we follow traditional SA (Qi et al., 2023a; b) to adopt multiple metrics. (1) Concordance index (CI), as a prevalent metric in SA, assesses models’ discrimination power. (2) Mean absolute error (MAE) gives the absolute error of model’s prediction on time-to-event. (3) Distribution calibration (D-cal(Haider et al., 2020) is a statistical test to evaluate models’ ability in calibrating survival distribution prediction; we mainly check if there is no significant difference in distribution between ground truth and prediction ($p>$ 0.05). For any method, we adopt 5-fold cross-validation to evaluate its performance and report the average on 5 folds. Please refer to Appendix C.4 for more details on evaluation metrics.
@@ 4.3 Analysis of Language-Encoded Prognostic Priors
Figure 2: Interpreting the survival prediction of VLSA via language-encoded prognostic priors. Top row shows language descriptions (simplified for better view) about prognostic visual features in WSIs. Detailed texts are provided in Appendix C.3. Middle row gives the most representative patches corresponding to each prognostic text. Last row presents each prognostic prior’s contribution to risk. We mainly examine the top three language priors in terms of contribution. The three examples are from the first three datasets. More results are shown in Appendix D.2.
Visualization We conduct a case study to further examine the role of language-encoded prognostic priors in our VLSA. As shown in Figure 2, we observe that valuable visual features can be highlighted by prognostic texts in most cases. These visual features, e.g., “differentiation of tumor cells” and “high cellular density”, could often provide important clues for cancer prognosis. We note that, such desirable behavior is largely attributed to the intrinsic property of VLMs, i.e., VLMs offer a well-aligned vision-language embedding space to enable us to manipulate visual features with language prior. We believe that this property would be particularly important to the MIL in which a model is usually required to aggregate numerous instances under weak supervision.
@@ 4.4 Analysis of Ordinal Inductive Biases
Ablation Study We conduct ablation studies to verify the effectiveness of our two ordinal inductive bias terms when the language-encoded prognostic priors ${\bm{P}}$ are used or not used in VLSA. Its results are shown in Table 3. We have two main observations from these results. (1) Ordinal survival prompt: when ${\bm{P}}$ is not used in VLSA, the improvements by imposing ordinality on survival prompts are 1.2% and 1.3% in average CI; when ${\bm{P}}$ is used, the improvements are relatively large (3.4% and 3.6%) on BRCA and are often marginal on the other four datasets. (2) Ordinal probability: also called ordinal IF, it always shows consistent improvements in terms of average MAE, ranging from 1.2 to 1.7, when ${\bm{P}}$ is used in VLSA or not. More discussions on the two inductive bias terms are provided in Appendix B.
@@
- w/o language-encoded prognostic priors
0.5971 0.5994 0.7853 0.5750 0.6840 0.6482 28.70 5
($\pm$ 0.033) ($\pm$ 0.086) ($\pm$ 0.015) ($\pm$ 0.064) ($\pm$ 0.070)
0.6037 0.6202 0.7893 0.6053 0.6836 0.6604 28.01 5
($\pm$ 0.043) ($\pm$ 0.046) ($\pm$ 0.018) ($\pm$ 0.065) ($\pm$ 0.036)
0.5997 0.6049 0.7846 0.5818 0.6769 0.6496 26.99 3
($\pm$ 0.033) ($\pm$ 0.104) ($\pm$ 0.016) ($\pm$ 0.056) ($\pm$ 0.065)
0.6083 0.6180 0.7908 0.6048 0.6908
- w/ language-encoded prognostic priors
@@
Discussion We note that there are two possible reasons for the marginal CI gain of two ordinal inductive biases when ${\bm{P}}$ is used in VLSA. (1) Ordinal incidence function, often helps to improve MAE, not CI, observed from the results of Table 3. (2) The effect of the ordinal inductive bias imposed on survival prompts may be pared down, when the learned visual features become more discriminative (due to the use of ${\bm{P}}$). This is because our survival prompts are learnable and they can be tuned better in optimization to align with the discriminative visual features. As a result, the effect of the ordinal bias term for prompts could be alleviated. We note that, although there is a marginal gain in CI when ${\bm{P}}$ is present, the effect of ordinal inductive biases on VLSA is often notable in the other terms, such as MAE and the ordinal Acc in Figure 3(a).
Figure 3: Ordinality visualization. (a) Heatmap of the similarity between any two learned survival prompts. Its horizontal axis places the first prompt to the last prompt from left to right; its vertical one does so from top to bottom. Acc is the accuracy of prompt ranking. The results are from the first three datasets. More results are shown in Appendix D.2. (b) Predictive probability (Proba.) w/o and w/ ordinality. The patients from test set are used for prediction. $\Delta$EMD is equal to $\text{EMD}({\bm{y}},\hat{{\bm{y}}}_{\text{w/o}})-\text{EMD}({\bm{y}},\hat{{\bm{y}}}_{\text{w/}})$. Vertical dashed line indicates individual time-to-event.
Ordinality Visualization (1) Survival prompts: we calculate the similarity between any two survival prompts learned by the model and show the similarity in heatmap. By comparing the heatmaps from the VLSA model with and without ordinal prompts, the effect of ordinal prompts can be examined qualitatively. As shown in Figure 3(a), we can see that survival prompts are better in ordinality when considering an ordinal inductive bias for them. (2) Incidence function: the case study shown in Figure 3(b) suggest that the ordinality constraint imposed on IF could help to decrease the EMD between $\hat{{\bm{y}}}$ and the ground truth. This often leads to the gains in MAE, as seen from Table 3.
4.5 Analysis of Data Efficiency (Few-shot Survival Analysis)
Here we focus on a few-shot learning (FSL) scenario in which only a few samples are available for training. We use this scenario to simulate the case of scarce training data and assess the performance of different SA methods in terms of data efficiency. Since existing works on SA rarely study FSL, this experiment follows common few-shot settings. Concretely, we randomly sample $s$ patients from each time-discrete class. For censored patients, their true class labels are partially given, so we adopt the KM method (Kaplan & Meier, 1958) to estimate them in advance for few-shot sampling. Moreover, due to small patient numbers, we run each single experiment 5 times and report the median metrics, following CONCH (Lu et al., 2024). MI-FewShot (Xu et al., 2024) is a non-parametric method designed for FSL, so we adopt it as an additional baseline for comparisons.
Table 4: Few-shot survival analysis for data efficiency evaluation.
@@ 4.5 Analysis of Data Efficiency (Few-shot Survival Analysis)
The results are shown in Table 4. (1) From these results, we find that our VLSA still often obtains the best performance in FSL scenarios. This indicates the data efficiency of our VLSA models. (2) When using only 16 shots, our VLSA can obtain comparable performance with nearly-all full-shot baseline models. (3) Furthermore, we observe that, in standard full-shot setting, two VL-based baselines, i.e., CoOp and OrdinalCLIP with ABMIL as their MIL aggregator, only show an improvement of 0.6% and 1.8% over vision-only ABMIL, respectively. By contrast, the average improvements seen in few-shot settings are around 6.5%, much larger than those in full-shot. This result also suggests the advantage of VL-based schemes over vision-only ones in data efficiency.